OpenStudy (anonymous):

find the maximum area of a rectangle whose perimrter is 30

OpenStudy (anonymous):

7.5 * 7.5 = 56.25 sq. units

OpenStudy (anonymous):

7.5*7.5=56.25sq.units

OpenStudy (anonymous):

7.5*7.5=56.25 sq units